# TSMixer: Lightweight MLP-Mixer Model for Multivariate Time Series Forecasting

### v1

@@ Abstract.
TSMixer outperforms state-of-the-art MLP and Transformer models in forecasting by a considerable margin of 8-60%. It also outperforms the latest strong benchmarks of Patch-Transformer models (by 1-2%) with a significant reduction in memory and runtime (2-3X).
Accepted in the Proceedings of the 29th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’23), Research Track, August 6–10, 2023, Long Beach, CA, USA
@@ 2. Related Work
MLP based sequence / time series models: Similar efforts have recently been put in the time series domain. Zhang et. al. (Zhang et al., 2022) proposed LightTS that is built upon MLP and two sophisticated downsampling strategies for improved forecasting. The DLinear model (Zeng et al., 2022a) also employs a simple linear model after decomposing the time series into trend and seasonality components. DLinear questions the effectiveness of Transformers as it easily beats Transformer-based SOTAs. Li et. al. (Li et al., 2022) proposed MLP4Rec, a pure MLP-based architecture for sequential recommendation task, to capture correlations across time, channel, and feature dimensions.
@@ 3.3.7. Forecast online reconciliation
Online hierarchical patch reconciliation head:444This “reconciliation” is different from the reconciliation in hierarchical forecasting (Hyndman and Athanasopoulos, 2021). Here, hierarchy is defined as an aggregation over a patch, and the reconciliation is done online during training..
@@ 4.2.2. Computational Improvements
PatchTST is considered the most compute-effective model across all time series Transformers, as patching drastically reduces the input size by a factor of s (stride) (Nie et al., 2022). In contrast, TSMixer not only enables patching but also completely eliminates the self-attention blocks. Hence, TSMixer shows significant computation improvement over PatchTST and other Transformer models. In Table 3, we highlight the speed-up and memory comparison of TSMixer over PatchTST. For analysis, we capture the following metrics: (i) Multiply-Add cumulative operations on the entire data per epoch (MACs), (ii) Number of model parameters (NPARAMS), (iii) Single EPOCH TIME and (iv) Peak GPU memory reached during a training run (MAX MEMORY). For this experiment, we trained TSMixer and PatchTST models in a single GPU node with the same hardware configuration in a non-distributed manner to report the results. To ensure fairness in comparison, we use the exact model parameters of PatchTST and TSMixer which were used in the error metric comparison reported in Table 2. In Table 3, we capture the average improvement of TSMixer over PatchTST for each performance metric across the three larger datasets (Electricity, Weather, Traffic) with =96. Since CI-TSMixer is purely MLP based, it significantly reduces the average MACs (by 4X), NPARAMS & MAX MEMORY (by 3X), and EPOCH TIME 555MACs and EPOCH TIME are highly correlated and in general create a similar relative impact across models. However, since PatchTST involves high parallelism across the attention heads, we observe a different relative impact w.r.t. MACs and EPOCH TIME. (by 2X). Even after enabling gated attention and hierarchy reconciliation, CI-TSMixer(G,H) still shows a good reduction in MACs & NPARAMS (by 3X), and EPOCH TIME & MAX MEMORY (by 2X). It is important to note that, when cross-channel reconciliation is enabled [i.e CI-TSMixer(G,H,CC)], the number of parameters becomes very high in TSMixer as compared to PatchTST. The reason is that the number of parameters in the cross-channel reconciliation head is dependent on the number of channels in the dataset that leads to this scaling effect. For example, since Electricity and Traffic datasets have a very high number of channels (i.e. 321 and 862 respectively), the number of parameters of the model also scales up, whereas weather (with only 21 channels) did not encounter any such scaling effect. Even PatchTST should show a similar scaling effect if the channel correlation is enabled in it. However, even with increased parameters, CI-TSMixer(G,H,CC) still shows a notable reduction in MACs (by 3X) and EPOCH TIME & MAX MEMORY (by 2X). The reason is that parameter scaling affects only the reconciliation head and not the backbone which primarily constitutes the total training time and memory. Thus, TSMixer and its variants can easily produce improved results over PatchTST in significantly less training time and memory utilization.
@@ A.1. Datasets
We use popular multivariate datasets provided in (Nie et al., 2022) for forecasting and representation learning. Weather666https://www.bgc-jena.mpg.de/wetter/ dataset collects 21 meteorological indicators such as humidity and air temperature. Traffic777https://pems.dot.ca.gov/ dataset reports the road occupancy rates from different sensors on San Francisco freeways. Electricity888https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 captures the hourly electricity consumption of 321 customers. ETT999https://github.com/zhouhaoyi/ETDataset (Electricity Transformer Temperature) datasets report sensor details from two electric Transformers in different resolutions (15 minutes and 1 hour). Thus, in total we have 4 ETT datasets: ETTM1, ETTM2, ETTH1, and ETTH2.
@@ A.2. Supplementary Details
TSMixer follows (Smith and Topin, 2017) to determine the optimal learning rate during training for better convergence. Early stopping with patience 10 is applied during training. TSMixer library is built using PyTorch and multi-GPU node training is enabled via Pytorch DDP101010. TSMixer can easily scale and cater to the needs of large-scale forecasting and can be deployed across multiple nodes in the Kubernetes clusters.

### v4

@@ Abstract.
TSMixer outperforms state-of-the-art MLP and Transformer models in forecasting by a considerable margin of 8-60%. It also outperforms the latest strong benchmarks of Patch-Transformer models (by 1-2%) with a significant reduction in memory and runtime (2-3X). The source code of our model is officially released as PatchTSMixer in the HuggingFace [Model]
[Examples].
Accepted in the Proceedings of the 29th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’23), Research Track. Delayed release in arXiv to comply with the conference policies on the double-blind review process. This work has been submitted to the KDD peer-review process on Feb, 02, 2023 (kdd, 2023).
@@ 2. Related Work
MLP based sequence / time series models:444Comparison with (Chen et al., 2023) is not included, as it was released in arXiv one month after the start of the KDD 2023 peer-review process (kdd, 2023), where our current paper has been submitted and got accepted. Similar efforts have recently been put in the time series domain. Zhang et. al. (Zhang et al., 2022) proposed LightTS that is built upon MLP and two sophisticated downsampling strategies for improved forecasting. The DLinear model (Zeng et al., 2022a) also employs a simple linear model after decomposing the time series into trend and seasonality components. DLinear questions the effectiveness of Transformers as it easily beats Transformer-based SOTAs. Li et. al. (Li et al., 2022) proposed MLP4Rec, a pure MLP-based architecture for sequential recommendation task, to capture correlations across time, channel, and feature dimensions.
@@ 3.3.7. Forecast online reconciliation
Online hierarchical patch reconciliation head:555This “reconciliation” is different from the reconciliation in hierarchical forecasting (Hyndman and Athanasopoulos, 2021). Here, hierarchy is defined as an aggregation over a patch, and the reconciliation is done online during training..
@@ 4.2.2. Computational Improvements
PatchTST is considered the most compute-effective model across all time series Transformers, as patching drastically reduces the input size by a factor of s (stride) (Nie et al., 2022). In contrast, TSMixer not only enables patching but also completely eliminates the self-attention blocks. Hence, TSMixer shows significant computation improvement over PatchTST and other Transformer models. In Table 3, we highlight the speed-up and memory comparison of TSMixer over PatchTST. For analysis, we capture the following metrics: (i) Multiply-Add cumulative operations on the entire data per epoch (MACs), (ii) Number of model parameters (NPARAMS), (iii) Single EPOCH TIME and (iv) Peak GPU memory reached during a training run (MAX MEMORY). For this experiment, we trained TSMixer and PatchTST models in a single GPU node with the same hardware configuration in a non-distributed manner to report the results. To ensure fairness in comparison, we use the exact model parameters of PatchTST and TSMixer which were used in the error metric comparison reported in Table 2. In Table 3, we capture the average improvement of TSMixer over PatchTST for each performance metric across the three larger datasets (Electricity, Weather, Traffic) with =96. Since CI-TSMixer is purely MLP based, it significantly reduces the average MACs (by 4X), NPARAMS & MAX MEMORY (by 3X), and EPOCH TIME 666MACs and EPOCH TIME are highly correlated and in general create a similar relative impact across models. However, since PatchTST involves high parallelism across the attention heads, we observe a different relative impact w.r.t. MACs and EPOCH TIME. (by 2X). Even after enabling gated attention and hierarchy reconciliation, CI-TSMixer(G,H) still shows a good reduction in MACs & NPARAMS (by 3X), and EPOCH TIME & MAX MEMORY (by 2X). It is important to note that, when cross-channel reconciliation is enabled [i.e CI-TSMixer(G,H,CC)], the number of parameters becomes very high in TSMixer as compared to PatchTST. The reason is that the number of parameters in the cross-channel reconciliation head is dependent on the number of channels in the dataset that leads to this scaling effect. For example, since Electricity and Traffic datasets have a very high number of channels (i.e. 321 and 862 respectively), the number of parameters of the model also scales up, whereas weather (with only 21 channels) did not encounter any such scaling effect. Even PatchTST should show a similar scaling effect if the channel correlation is enabled in it. However, even with increased parameters, CI-TSMixer(G,H,CC) still shows a notable reduction in MACs (by 3X) and EPOCH TIME & MAX MEMORY (by 2X). The reason is that parameter scaling affects only the reconciliation head and not the backbone which primarily constitutes the total training time and memory. Thus, TSMixer and its variants can easily produce improved results over PatchTST in significantly less training time and memory utilization.
@@ A.1. Datasets
We use popular multivariate datasets provided in (Nie et al., 2022) for forecasting and representation learning. Weather777https://www.bgc-jena.mpg.de/wetter/ dataset collects 21 meteorological indicators such as humidity and air temperature. Traffic888https://pems.dot.ca.gov/ dataset reports the road occupancy rates from different sensors on San Francisco freeways. Electricity999https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 captures the hourly electricity consumption of 321 customers. ETT101010https://github.com/zhouhaoyi/ETDataset (Electricity Transformer Temperature) datasets report sensor details from two electric Transformers in different resolutions (15 minutes and 1 hour). Thus, in total we have 4 ETT datasets: ETTM1, ETTM2, ETTH1, and ETTH2.
@@ A.2. Supplementary Details
TSMixer follows (Smith and Topin, 2017) to determine the optimal learning rate during training for better convergence. Early stopping with patience 10 is applied during training. TSMixer library is built using PyTorch and multi-GPU node training is enabled via Pytorch DDP111111. TSMixer can easily scale and cater to the needs of large-scale forecasting and can be deployed across multiple nodes in the Kubernetes clusters.